OpenStudy (anonymous):

-4a+12+8a=7a-9a Do i just combine the like terms first? Or am i supposed to try and get the A by itself?

OpenStudy (midhun.madhu1987):

yeah... combine the like terms... and simplify

OpenStudy (anonymous):

So would it be 4a+12=7a-9a

OpenStudy (anonymous):

4a+12=12a

OpenStudy (anonymous):

4a+12=-2 I meant

OpenStudy (midhun.madhu1987):

 find a.